.254 mm into inches

.254 mm into inches
Answer #1

Millimeters (mm) is the unit of the length in matric system

1 inch = 25.4mm

Now we have to convert 0.254mm to inches

So

.254mm= x ( let x be the inch we have to find )

25.4/100= x ( now we know 25.4 is equal 1 inch )

1 (inch) /100=x

0.01 inch =x

Hence x is equal to 0.01 inch

So 0.254 is equal to 0.01 inch

Hope you find this helpful thank you ☺️